the atomic weight of boron is 10.811. the natural abundance of 10B is 19.91%. Determine the atomic weight of the only other natural isotope of boron
1
Expert's answer
2019-09-09T04:50:27-0400

Another natural boron isotope is xB. xB and 10B form the average atomic weight of boron:


"10 * 0.1991 + X * (1-0.1991) = 10.811", where X -- the atomic mass of another boron isotope;


"1.991 + 0.8009*X = 10.811"


"0.8009*X = 8.82"


"X = 11"


The atomic weight of the only other natural isotope of boron is 11.